OpenStudy (lynfran):

Find the distance from (0,0) to the line 4x-4y=9

OpenStudy (lynfran):

ok so i work this problem ..but keep getting a negative distance...can that be??

Vocaloid (vocaloid):

nope, distance is always positive

OpenStudy (lynfran):

i know thats what bothers mee

OpenStudy (anonymous):

probably be easier if you write the line as \[y=x-\frac{9}{4}\]

OpenStudy (lynfran):

but dont we have to use the normal/ perpendicular formula \[d=\frac{ Ax+By+C }{ \pm \sqrt{A^2+B^2} }\]

OpenStudy (anonymous):

oh okay not familiar with that what are x and y in that formula ?

OpenStudy (lynfran):

the point given (0,0)..

OpenStudy (anonymous):

OOH

OpenStudy (anonymous):

then this gotta be simple right?

OpenStudy (lynfran):

but i keep getting a negative distance

OpenStudy (anonymous):

\[\frac{9}{\sqrt{4^2+4^2}}\]

OpenStudy (anonymous):

the numerator should be the absolute value your formula has a silly \(\pm\) in it which just means make sure it is positive

OpenStudy (lynfran):

so 1.5909 would be the distance??

OpenStudy (anonymous):

\[d=\frac{ Ax+By+C }{ \pm \sqrt{A^2+B^2} }\] should be \[d=\frac{ |Ax_0+By_0+C| }{ \sqrt{A^2+B^2} }\]

OpenStudy (anonymous):

idk are you supposed to give a decimal approximation, or the answer as a radical?

OpenStudy (lynfran):

i didnt say

OpenStudy (lynfran):

ok i have another question Determine K so that the distance from the origin to the line y=km+9; p=6 will be stated

OpenStudy (lynfran):

@sleepyhead314 .. wat do u think?

OpenStudy (sleepyhead314):

well for your new question, if you plug in what they gave into the equation... \[\frac{ 9 }{ \sqrt{k^2 + 1} } = 6\]

OpenStudy (lynfran):

where did u get that equation frm..?

OpenStudy (lynfran):

did u derive that ?

OpenStudy (sleepyhead314):

ehh? the equation you gave before...? \[\frac{ Ax + By + C}{ \sqrt{A^2 + B^2} }\]

OpenStudy (lynfran):

o its the same so 9=C but what about A and B?

OpenStudy (sleepyhead314):

I just plugged in A = -k B = 1 C = 9 x = 0 y = 0 because y = km + 9 --> -km + y = 9

OpenStudy (lynfran):

ok so we now solve for K?

OpenStudy (sleepyhead314):

yep :)

OpenStudy (lynfran):

\[9=6(\sqrt{K^2+1})\]\[\frac{ 3 }{ 2}=\sqrt{k^2+1}\]\[\frac{ 9 }{ 4 }=k^2+1\]\[\frac{ 5 }{ 4 }=K^2\]\[\frac{ \sqrt{5} }{ 2 }=k\]

OpenStudy (sleepyhead314):

that's what I got :)

OpenStudy (lynfran):

ok is there a way to check if its correct though?

OpenStudy (lynfran):

like when we solve equation we can sub the value we find back into the equation to see that its correct..eg 4x+2y=0 so y=0 and x=0 4(0)+2(0)=0 like this...^^

OpenStudy (sleepyhead314):

I suppose you could plug in A = (sqrt(5)/2) B = 1 C = 9 x = 0 y = 0 and make sure you get 6 ?

OpenStudy (lynfran):

ok cool.. thanks to all